A stone dropped from the top of the tower travel 25m in the last second of its motion. If g is 10m/s, height of the tower is?

Grade:12th

1 Answers

Kruthik
6 years ago
The answer is 45metres
Let’s consider the last second of its journey.
We have S=25m, a=g, t=1s
Thus, by using s=ut + 1/2at2, we get u=20ms-1
 Now, during the time before the last second, let h be the distance covered. 
Here, S=h, a=g, v=20ms-1,u=0
Therefore, by using V2=U+2aS, we have S=h=20m.
Now, the total height is h+25 = 45meters. 
Hope that helped you. ✌️